### v1

@@ 6 Conclusions
We have reported a pair of very dissimilar shocks observed by Solar Orbiter at 0.83 au from the Sun. The first was a quasi-parallel shock which satisfied all expectations, and showed strong non-equilibrium plasma signatures as seen in fig. 5. The second, quasi-perpendicular shock exhibited unusual, rapid ion dynamics which were nevertheless well resolved with the EPD sensors aboard Solar Orbiter. Many of the variations of the energetic-particle fluxes were observed to be at the ion cyclotron time scale or even more rapid. The high time and energy resolution offered by Solar Orbiter’s in situ instruments is critical for understanding the micro-physics of particle dynamics and acceleration at traveling interplanetary shocks. The rapid variations and dispersive features would not have been visible with the lower resolution provided by similar instruments on previous missions. They would probably have been interpreted as broad diffusive signatures.
@@ Acknowledgements.
Appendix A STEP measurements during time period 1)

### v2

@@ 6 Conclusions
It was the best of times, it was the worst of times, …- for the thermal and suprathermal particle populations in the vicinity of this pair of very dissimilar shocks observed by Solar Orbiter at 0.83 au from the Sun. The first shock was a quasi-parallel shock and showed strong non-equilibrium plasma signatures as seen in fig. 5. Because it ran into a small ICME with its low level of IMF fluctuations, suprathermal particles were not reflected back into the shock which therefore did not have an appreciable effect on energetic particles. The second, quasi-perpendicular shock exhibited unusual, rapid ion dynamics which were nevertheless well resolved with the EPD sensors aboard Solar Orbiter. Many of the variations of the energetic-particle fluxes were observed to be at the ion cyclotron time scale or even more rapid. The high time and energy resolution offered by Solar Orbiter’s in situ instruments is critical for understanding the micro-physics of particle dynamics and acceleration at traveling interplanetary shocks. The rapid variations and dispersive features would not have been visible with the lower resolution provided by similar instruments on previous missions. They would probably have been interpreted as broad diffusive signatures.
@@ Acknowledgements.
Appendix A Determination of the critical Mach number
The concept of shock criticality is well–defined when shocks are considered thorough a fluid approach, where a critical Mach numbers arises when the downstream flow speed equals the downstream sound speed and a shock solution is only possible when some other dissipation process is introduced (Burgess & Scholer 2015). In the collisionless case, such process is mediated by ion reflection, and the definition of the critical Mach number becomes somewhat vague, motivating several studies investigating supercriticality in different contexts (e.g., Amano & Hoshino 2010; Ha et al. 2018). The critical Mach number is determined by three quantities, the upstream $\langle\theta_{Bn}\rangle$, plasma $\beta$, and the ratio of specific heats, $\gamma$, of the solar wind plasma Edmiston & Kennel (1984). The upstream values of the first two quantities for the two shocks considered in this paper are given in tab. 1. The value of the third quantity, $\gamma$, cannot be measured directly and requires further consideration. The solar wind consists primarily of protons and doubly-ionized helium ions with a small admixture of heavy ions which, together with electrons, all contribute to the solar wind ratio of specific heats. de Avillez et al. (2018) investigated a plasma with solar composition and found that $\gamma$ is not constant, but depends on composition, temperature, and even the exact shape of the velocity distribution functions of the plasma constituents. The solar wind is depleted in helium Geiss (1982); von Steiger et al. (2000) compared to the Sun, and is further depleted in the slow solar wind in which the two shocks were embedded Bochsler (2007). Using a lower limit of $v_{\rm th}=10$ km/s for the thermal speed of protons (see the corresponding panels in figs. 4 and 7) we arrive at $T>10^{5}$ K. Figure 4 of de Avillez et al. (2018) shows clearl that we can safely assume $\gamma\approx 5/3$ for these solar wind conditions. Using this value, we determined the critical Mach number, $M_{C}=1.7$ ($M_{C}=2.2$), for shock #1 (#2) by interpolating in fig. 4 of Edmiston & Kennel (1984). Thus we independently confirmed our conclusion that they were both supercritical based on the presence of the foot in both shocks Bavassano-Cattaneo et al. (1986).
Appendix B STEP measurements during time period 1)